can you please show to me the proof of the sum of the exterior angles of any polygons which is equal to 720 degrees..thank you.
Answer by Theo(13342)

The sum of the exterior angles of a polygon is equal to 360 degrees (not 720).
let I = the interior angle of a polygon.
let E = the exterior angle of a polygon.
The exterior angle of a polygon is equal to 180 - the interior angle of a polygon.
This is expressed as:
E = 180 - I
The sum of the interior angles of a polygon is given by the formula:
sum(I) = (n-2)*180 where n is the number of sides of the polygon.
From this formula, the interior angle of a formula is calculated as:
I = (n-2)*180/n
Since the exterior angle of a polygon is always supplementary to the interior angle of a polygon, this means that:
E = 180 - (n-2)*180/n
simplify this formula to get:
E = 180 - (180n-360)/n
Since 180 = 180n/n, this equation can be rewritten as:
E = 180n/n - (180n-360)/n
This can be further simplified to:
E = 180n/n - 180n/n + 360/n
Combine like terms and you get:
E = 360/n
Multiply both sides of this equation by n to get:
n*E = 360
Since n*E is equal to the sum of the exterior angles of any polygon, this means that:
The sum of the exterior angles of any polygon is equal to 360 degrees.